OpenStudy (anonymous):

Aspartame is an artificial sweetener that is 160 times sweeter than sucrose (table sugar) when dissolved in water. It is marketed as Nutra-Sweet. The molecular formula of aspartame is C14H18N2O5. a- Calculate the molar mass of aspartame. b- what amount (moles) of molecules are present in 10.0 g aspartame? c-Calculate the mass in grams of 1.56 mol aspartame. d-What number of molecules are in 5.0 mg aspartame? e-What number of atoms of nitrogen are in 1.2g aspartame?

OpenStudy (aaronq):

a) use the periodic table b) Use: \(\sf moles=\dfrac{mass}{molar~mass}\) c) use the formula from b) d) use the formula from b) and \(\sf moles=\dfrac{particles}{avogadro's~number }\) (in this case, molecules are the particles) e) Find the mass percent composition of N in aspartame: \(\sf mass~percent~comp=\dfrac{N's~molar~mass}{Molar~mass~of~aspartame}*100\%\) multiply the mass of aspartame (1.2 g) by the mass percent composition of N then (following the formula from b) multiply it by avogadro's number

OpenStudy (anonymous):

thank you soooo much @aaronq

OpenStudy (aaronq):

no problems